Circular No. 9056

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2009gj IN NGC 134 C. J. Stockdale, B. Rentz, and C. M. Vandrevala, Marquette Uni-

versity; K. W. Weiler, Naval Research Laboratory; S. Immler, Goddard Space Flight Center; S. D. Van Dyk, Spitzer Science Center, California Institute of Technology; N. Panagia, Space Telescope Science Institute and Observatory of Catania; J. M. Marcaide, University of Valencia; D. Pooley, University of Wisconsin at Madison; R. A. Sramek, National Radio Astronomy Observatory; and S. Ryder, Anglo-Australian Observatory, report the detection of radio emission near the position of the type-IIb supernova 2009gj (CBETs 1856, 1858) with the Very Large Array radio telescope. Flux density measurements were made at 8.46 GHz (3.5 cm) of 0.33 +/- 0.04 mJy on June 27.43 UT, 0.59 +/- 0.04 mJy on July 2.56, and 1.02 +/- 0.04 on July 11.51. The measured position of the radio emission of α = 0^{h}30^{m}28^{s}.54, δ = −33^{o}12′56′′.2 (equinox 2000.0) is in good agreement with the measured optical position end figures 28^{s}.56, 56′′.0 (CBET 1856). The following 3σ upper limits were measured: at 22.46 GHz (1.3 cm), 1.9 mJy on June 28.45 and 1.5 mJy on July 5.51; at 4.86 GHz (6.0 cm), 0.19 mJy on June 29.44 and 0.38 mJy on July 2.60; and at 1.43 GHz (20 cm), 2.32 mJy on July 2.64. The radio emission has likely been caught on the rise at 3.6 cm. If the object behaves like other known type-IIb supernovae, it should soon start increasing rapidly at longer radio wavelengths. Further study in all wavelength bands, especially radio and x-ray, is warranted. Radio observations at the Very Large Array are continuing.

COMETS C/2009 C5 AND C/2009 D2–D4 (SOHO) Additional Kreutz sungrazers have been found on SOHO website im-

ages (cf. IAUC 9055; ML = M. Kaluzny). C/2009 C5 was faint and slightly diffuse, peaking at mag ∼ 7. C/2009 D2 and C/2009 D3 were small and stellar in appearance, peaking at mag ∼ 7 and 7–7.5, respectively. C/2009 D4 showed a long, thin tail that was visible in C3 images (which showed a small head that peaked at mag ∼ 2–3 and a tail length of ∼ 1^{o}) as well as in C2 SOHO images and in SECCHI COR2-A and COR2-B images.

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 C5 Feb. 11.529 21^{h}54.^{m}9 −14^{o}33′ C3  ML 2009-M54
C/2009 D2  19.179 22 27.5 −12 11 C3  JR 2009-M54
C/2009 D3  19.321 22 26.9 −11 52 C3  BZ 2009-M54
C/2009 D4  21.943 22 53.8 −11 59 C3/2* TH 2009-M55
2009 July 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green